# UDC: A Unified Neural Divide-and-Conquer Framework for Large-Scale Combinatorial Optimization Problems

### v1

@@ A.1 Neural Combinatorial Optimization
Over the past few years, extensive research on learning-based methods has been applied to various CO problems. Pointer Network [55] first applies the deep learning technique to solve NP-hard CO problems, after that, [28] designs an RL-based framework for training NCO solvers. In recent years, more diverse and sophisticated NCO frameworks have been proposed [56, 57]. Among them, the mainstream methods include (RL-based) constructive solvers [50], (SL-based) sub-path solvers [8], the heatmap-based solvers [35], and the neural divide-and-conquer methods [13]. These categories of methods demonstrate diverse advantages across different scales and CO problems.
@@ A.2 Constructive Neural Solvers
The basic idea of constructive neural solvers has been illustrated in Section 2.2. Up to now, the constructive neural solver is the most popular method for solving CO problems. Introducing the self-attention mechanism [29], the Attention Model [50] demonstrates the applicability of various VRPs and the efficiency in solving small-scale VRP instances. AM achieves outstanding results in six VRPs including TSP, CVRP, SDVRP, PCTSP, SPCTSP, and OP. Recent research mainly focuses on the two representative CO problems, TSP and CVRP. Some studies propose more advanced problem-solving strategies [58], others introduce more efficient network architectures [9, 51], and some focus on training by upgrading the calculation of the loss function [5, 59] to better escape potential local optima.
By redesigning the feasibility mask in each autoregressive construction step, constructive solvers can be effectively applied to most CO problems. Therefore, although many new model structures have been proposed for different CO problems (such as [41] for the Pickup and Delivery Problem (PDP)), ordinary frameworks are usually still feasible. This article adopts the model architectures proposed in POMO [5] and ICAM [16] for most CO problems for convenience. However, for two special problems, MIS and min-max mTSP, conventional constructive solvers are unable to process constraints in their encoders. Therefore, we specifically use one-shot AGNN (consistent with the dividing policy) in sub-MIS and DPN [60] to solve sub-min-max mTSP.
@@ A.3 Heatmap-based Solvers
Recent heatmap-based solvers propose methods to improve the performance of GNN-based frameworks [31, 32]. Attn-MCTS [37] proposes a heuristic method to merge the smaller heatmap obtained from small graphs into a larger heatmap of TSP. More importantly, Attn-MCTS also proposes an effective Monte-Carlo tree search method (MCTS) based on large-scale heatmaps. Therefore, afterward works turn to pay more attention to providing better initial heatmaps for the following Monte Carlo tree search [61]. DIMES [35] proposes a meta-learning method based on reinforcement learning for stability in training, while DIFUSCO [39] and T2T [38] use supervised learning to learn a diffusion process [62] of heatmap generation.
@@ A.4 Neural Divide-and-Conquer Methods
The Table 1 in the main paper offers a brief overview of all the existing neural divide-and-conquer methods. These methods and the proposed UDC follow the same two-stage solution generation framework. Different from the total neural framework of UDC, some neural divide-and-conquer methods strategically incorporate heuristics in either the dividing or conquering stages to enhance performance or efficiency [22, 23, 21, 13], which is hard to generalize to other CO problems.
@@
LCP [24], H-TSP [20], RBG [23], TAM [12], and GLOP [13] (specifically for CVRP and PCTSP) can use only neural networks in both stages, which enhances the applicability of methods to handle CO problems without expert experience. However, these methods face challenges in terms of effectiveness. We believe the deficiency in performance is highly related to the shortcomings in their chosen training schemes. Among them, LCP [24] adopts a totally separate training scheme, employing a constructive neural solver (seeder) to generate diversified initial solutions (seeds) and another constructive neural solver (reviser) for sub-path solving. However, the seeders and revisers are never trained together and the objective functions of its seeders and revisers are not directly related to the objective function of the original CO problem (i.e., belonging to a separate training scheme), so its effectiveness largely depends on randomness.
@@
TAM [12], GLOP [13] (for CVRP and PCTSP), and RBG [23] adopt a similar training scheme, with only differences in the solver for the dividing policy (i.e., TAM uses a constructive neural solver, and GLOP chooses a heatmap-based solver). They pre-train a conquering policy on a preempted dataset and directly use them as local solvers (i.e., operators) for their initial solution. Such a training scheme fits the Pre-train + Train Dividing variant in Section 5.1 which cannot converge to a good performance. The pre-trained conquering policy is hard to ensure optimal so such a training scheme will probably lead to local optima. Moreover, as another drawback of TAM and GLOP, they decompose CVRP and PCTSP into TSP for convenience, but such a conquering strategy cannot convert sub-optimal initial partitions into optimal ones, affecting the accessibility of optimal solutions.
@@
H-TSP [20] focuses on the TSP problem and designs a lighter convolutional neural network and clustering mechanism for initial solution generation compared to LCP. It uses pre-trained sub-TSPs to achieve joint training but similar to LCP, it does not align the optimization objectives of the outer network with the overall objective function as well, which limits its effectiveness. In addition, Section 5.1 also demonstrates that the pre-training before a joint (i.e., unified) training stage will also lead to local optima and be harmful to convergence.
Appendix B Problem & Sub-problem Definition
@@
TSP. Traveling salesman problem (TSP) is one of the most representative COPs [63]. Let $\{D=d_{j,k},j=1,\ldots,N,k=1\ldots,N\}$ be the $N\times N$ cost metrics of a $N$-nodes TSP problem, where $d_{j,k}$ denotes the cost between nodes $k$ and $j$, the goal is to minimize the following objective function:
@@
where the solution $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{N})$ ($\tau=N$) is a permutation of all nodes. All the feasible solutions (i.e., in $\omega$ of Eq. (1)) satisfy the constraint of node degree being 2, and containing no loop with lengths less than N.
The sub-TSP in UDC is usually considered as an open-loop TSP problem [64] (i.e., SHPP in GLOP [13]). For the sub-cost metrics $\{D_{s}=d^{s}_{j,k},j=1,\ldots,n,k=1\ldots,n\}$, the objective function of sub-TSP solution $\boldsymbol{s}=(s_{1},s_{2},\ldots,s_{n})$ is as follow:
$\mbox{minimize}\quad f(\boldsymbol{s})=\sum_{t=1}^{n-1}d^{s}_{x_{t},x_{t+1}}$ (8)
The feasibility set of sub-TSP adds additional selection constraints for the starting and ending points (i.e., the first and last node in the original solution segments). UDC trains the TSP on various scales for every $n$ nodes, from $N$=500 to $N$=1,000 (i.e., for easier implement, sampling from $N=\{500,600,700,800,900,1,000\}$ when $n$=100). Since the nodes in each sub-problem always come from a compact region of the original instance, so we follow the normalization approach in GLOP and conduct a coordinate transformation to coordinates of sub-problems. The specific formula is detailed in Eq. (18).
CVRP. CVRP incorporates the vehicle capacity as a restriction while considering the total distance as its objective function. Each CVRP contains a depot and several customers, on a cost matrix $\{D=d_{j,k},j=0,\ldots,N,k=0\ldots,N\}$, the CVRP can be expressed as follows:
@@
$\sum_{i\in\boldsymbol{\rho}^{j}}\delta_{i}\leq C,\quad j=1,\ldots,q.$
where $\boldsymbol{x}$ is a solution representing the complete route of the vehicle and consists of $q$ sub-routes $\{\boldsymbol{\rho}^{1},\boldsymbol{\rho}^{2},\ldots,\boldsymbol{\rho}^{q}\}$. Each sub-route starts from the depot $x_{0}$ and backs to $x_{0}$, $n_{j}$ represents the number of customer nodes in it. $n=\sum_{j=1}^{q}n_{j}$ is the total number of customer nodes; $\delta_{i}$ denotes the demand of node $i$; $D$ denotes the capacity of the vehicle. For testing random CVRP instances, we follow the common setting in [13, 14], setting $C=100$ for CVRP500, $C=200$ for CVRP1,000, $C=300$ for CVRP2,000 and larger scale. In training, the capacity is randomly sampled from 100 to 200 [6, 16].
Similar to sub-TSP, Sub-CVRP also restricts the starting node and ending node. Moreover, sub-CVRP specially processes the capacity of the first and last routes to ensure the legality of the merged solution. As shown in the right half of Figure 4, the capacity of the first and last route is restricted to the rest capacity of the connected partial sub-route. Afterward, each two adjunct capacities (e.g., the pair of capacities shown in Figure 4) is normed to make the sum of 1. We followe the setting of BQ-NCO [14] in representing CVRP solutions, converting the solutions into an aligned solution flag which is set to 1 for nodes before a depot.
Figure 4: The process of sub-CVRP.
OP. The Orienteering Problem (OP) assigns each node a prize $\rho_{i}$ and the goal is to maximize the total prize of all collected nodes, while keeping the total length of the route below a maximum length $T$. OP contains a depot node as the starting node of the route, and we follow the general OP (distance) (i.e., OP(dist)) setting [50, 51, 59, 14], generating the value of $\rho_{i}$ based on its distance to the depot. We evaluate the UDC on OP500, OP1,000, and OP2,000 and set the $T$ on all scales to 4. The solution of OP is an $\tau$-length ordered sequence $\boldsymbol{x}=(x_{0},\ldots,x_{\tau})$ ($x_{0}$ represents be depot, $\rho_{0}$ = 0). Let $\{D=d_{j,k},j=0,\ldots,N,k=0\ldots,N\}$ be the $N\times N$ cost metrics of a $N$-nodes TSP problem, the objective function can be formulated as:
minimize $f(\boldsymbol{x})=\sum_{i\in\boldsymbol{x}}\rho_{i},$ (10)
@@
For generating OP instances, we follow the setting in AM [50], $\hat{\rho}_{i}=1+\lfloor 99\frac{d_{0,i}}{\text{max}_{j=1}^{N}d_{0j}}\rfloor$, $\rho_{i}=\frac{\hat{\rho_{i}}}{100}$. The sub-OP in UDC contains a segment of the original solution while also randomly introducing unselected nodes. By solving the sub-OP, the conquering stage has the potential to improve a low-quality initial solution to optimal. There is no normalization step in the conquering stage when addressing OP. Sub-OP restricts the total lengths of routes not to surpass the original total length instead of $T$. One of the $\lfloor\frac{N}{n}\rfloor$ sub-OPs contains the depot and the depot is restricted to be contained in the sub-solution of these sub-OPs.
In testing, solving sub-OP limits the path length not to exceed the original total length, so continuously executing the conquering stage will result in a gradually shorter total path length $l(\boldsymbol{x}_{r})$. So, we count the margin (i.e., $T-l(\boldsymbol{x})_{r}$) of the constraint $T$ in the current solution. In each conquering stage, we add the margin $T-l(\boldsymbol{x})_{r}$ to random one of the $\lfloor\frac{N}{n}\rfloor$ sub-OP problems.
PCTSP. With a similar optimization requirement compared to OP, PCTSP is also considered a typical CO problem [50]. PCTSP gives each node a prize $\rho_{i}$ and a penalty $\beta_{i}$. The goal is to minimize the total length of the tour plus the sum of penalties for nodes that are not visited while collecting at least a given minimum total prize. The objective function is defined as follows:
@@
We follow the settings in AM [50] and GLOP [13] to generate instances, $\rho_{i}\sim\text{Uniform}(0,\frac{4}{N})$, $\beta_{i}\sim\text{Uniform}(0,3\cdot\frac{K^{N}}{N})$. We follow the setting in [13], define $K^{N}=9,12,15$ for $n=500,1,000,2,000$.
Compared to OP, UDC uses a similar way to process sub-PCTSP, introducing random unselected nodes in solution segments and restricting the appearance of the depot. Specifically, sub-PCTSP normalizes the constraint of $\rho$ of each sub-PCTSP to 1. (The necessity of this procedure is discussed in Appendix E). Existing methods GLOP [13] divides the PCTSP into sub-TSP, such a problem-specific design not only harms the accessibility to optimal solutions once the node set of optimal solutions is inconsistent with the one of $\boldsymbol{x}_{0}$, but also harms the generation ability to general CO problems.
SPCTSP The outstanding results on SPCTSP evaluate the capability of the proposed UDC in dealing with uncertainty. In the SPCTSP, the expected node prize is known upfront, but the real collected prize only becomes known upon visitation. The expected node prize $\rho^{*}_{i}\in\text{Uniform}(0,2\rho_{i})$ is used to process sparse graph $\mathcal{G}_{D}$, heatmap $\mathcal{H}$, and the encodings of conquering policy, while the real prize $\rho_{i}$ is only used to process the current status and constraints when constructing solutions of SPCTSP and sub-SPCTSP.
KP. Besides VRPs, KP is also a typical CO problem which can be formulated as follows:
minimize $f(\mathbb{S})=-\sum_{i\in\mathbb{S}}v_{i},$ (12)
subject to $\sum_{i\in\mathbb{S}}w_{i}\leq W.$
where $\mathbb{S}\subseteq\{1,2,...,n\}$ represents the selection item list in a knapsack and $v_{i}\sim\text{Uniform}(0,1)$ and $w_{i}\sim\text{Uniform}(0,1)$ is the value and weight of candidate objects. In each conquering step, sub-KPs are randomly sampled. Sub-kp and KP are completely identical in form, with only distinction on the number of decision variables. In training, UDC disables the Reunion stage in the DCR training method for KP, and the capacity $W$ is sampled from 50 to 100. In testing, UDC adopts the same approach as OP in introducing the margin of maximum capacity ($W-\sum_{i\in\mathbb{S}}w_{i}$) to sub-KPs, adding the current margin (i.e., capacity minus current load) to a random sub-KP.
OVRP. As a variant on CVRP, OVRP only changes the objective function calculation as follows:
@@
ATSP. ATSP is a variant version of TSP with input being an asymmetric matrix. When dealing with sub-ATSP, different from sub-TSP, we employ a matrix-based special normalization method to normalize the longest item (i.e., edge) in the matrix to 1.
MIS. MIS is originally defined on a graph $\mathcal{G}=\{\mathbb{V},\mathbb{E}\}$ where $N=|\mathbb{V}|$ so $\mathcal{G}_{D}=\mathcal{G}$. The goal of MIS is to maximize the number of visited nodes and ensure that there are no edge connections $\boldsymbol{e}\in\mathbb{E}$ between two visited nodes. We follow the MIS instance generation process of ER[700-800] with p=0.15 [35]. The nodes in sub-MIS are randomly selected and nodes connecting to visited nodes are restricted unaccessible. The conquering policy for sub-MIS uses a Heatmap-based solver as well (i.e., AGNN in UDC) with non-autoregressive construction methods.
Min-max mTSP. As a variant of VRP, the min-max vehicle routing problem (min-max VRP), instead of minimizing the total length of all routes (i.e., min-sum), seeks to reduce the length of the longest one among all the routes (i.e., min-max). Min-max VRP also has important application value in real life [65, 66] and researchers also pay attention to solving them with deep learning techniques. Min-max mTSP restricts the number of total routes to $M$, the first node in $N$ nodes is the depot. Following the solution definition of CVRP (present in the form of a solution flag as well), the objective function is as follows:

### v2

@@ A.1 Neural Combinatorial Optimization
Over the past few years, extensive research on learning-based methods has been applied to various CO problems. Pointer Network [55] first applies the deep learning technique to solve NP-hard CO problems, after that, [28] designs an RL-based framework for training NCO solvers. In recent years, more diverse and sophisticated NCO frameworks have been proposed [56, 57]. Among them, the mainstream methods include (RL-based) constructive solvers [50], (RL-based) improvement-based solvers [58, 59], (SL-based) sub-path solvers [8], the heatmap-based solvers [35], and the neural divide-and-conquer methods [13]. These categories of methods demonstrate diverse advantages across different scales and CO problems.
@@ A.2 Constructive Neural Solvers
The basic idea of constructive neural solvers has been illustrated in Section 2.2. Up to now, constructive neural solvers are the most popular neural method for solving CO problems. By employing the self-attention mechanism [29], the Attention Model [50] demonstrates the applicability among various VRPs and the efficiency in solving small-scale VRP instances. AM achieves outstanding results in six VRPs including TSP, CVRP, Split Delivery Vehicle Routing (SDVRP), PCTSP, SPCTSP, and OP. Afterward, researchers mainly focus on the two most representative CO problems, TSP and CVRP. Some studies propose more advanced problem-solving strategies [60], others present more efficient network architectures [9, 51], and some focus on designing better loss functions [5, 61] to better escape potential local optima.
By employing a corresponding feasibility mask in each autoregressive construction step, constructive solvers can be effectively applied to most CO problems. Therefore, although many new model structures have been proposed for different CO problems (such as [62] for the Pickup and Delivery Problem (PDP)), ordinary frameworks are usually still feasible. This article adopts the model architectures proposed in POMO [5] and ICAM [16] for most CO problems for convenience. However, for some special CO problems, e.g., MIS, ATSP, and min-max mTSP, conventional constructive solvers are unable to process the problem-specific constraints in their encoders. Therefore, we specifically use one-shot AGNN (consistent with the dividing policy) in sub-MIS, MatNet [41] for sub-ATSP, and DPN [63] to solve sub-min-max mTSP.
@@ A.3 Heatmap-based Solvers
Recent heatmap-based solvers propose methods to improve the performance of GNN-based frameworks [31, 32]. Attn-MCTS [37] proposes a heuristic method to merge the smaller heatmap obtained from different regions of a TSP instance into a larger heatmap of TSP. More importantly, Attn-MCTS also integrates an effective Monte-Carlo tree search algorithm (MCTS) based on large-scale heatmaps. Therefore, the following heatmap-based works turn to pay more attention to providing better initial heatmaps for the following Monte Carlo tree search [64, 38]. DIMES [35] proposes a meta-learning method based on reinforcement learning for stability in training, while DIFUSCO [39] and T2T [65] use supervised learning to learn a diffusion process [66] of heatmap generation.
@@ A.4 Neural Divide-and-Conquer Methods
The Table 1 in the main part offers a brief overview of all the existing neural divide-and-conquer methods. These methods and the proposed UDC follow the same two-stage solution generation framework. Instead of being a fully neural method as UDC, some neural divide-and-conquer methods strategically incorporate problem-specific heuristics in either the dividing or conquering stages to enhance performance or efficiency [22, 23, 21, 13]. Such a setting undermines the possibility to generalize to other CO problems.
@@
LCP [24], H-TSP [20], RBG [23], TAM [12], and GLOP [13] (for CVRP and PCTSP) can use only neural networks in both stages, which enhances the applicability of methods to handle general CO problems without expert experience. However, these methods face challenges in terms of solution qualities. To tackle the instability in unified training [20], all these methods adopt a separate training scheme and we believe there are inherent shortcomings in the separate training schemes. Among them, LCP [24] employs a constructive neural solver (seeder) to generate diversified initial solutions (seeds) and another constructive neural solver (reviser) for sub-path solving. The seeders and revisers are trained separately and the objective functions of its seeders and revisers are not directly related to the objective function of the original CO problem (i.e., belonging to a separate training scheme), so its effectiveness largely depends on randomness.
@@
TAM [12], GLOP [13] (for CVRP and PCTSP), and RBG [23] adopt a similar training scheme, with only using different solvers for the dividing policy (i.e., TAM uses a constructive neural solver, and GLOP chooses a heatmap-based solver). These methods pre-train a conquering policy on a preempted dataset and directly use them as local solvers (i.e., operators) for their initial solution. Such a training scheme fits the Pre-train + Train Dividing variant in Section 5.1 which can hardly converge to a good performance. The pre-trained conquering policy will usually generate sub-optimal sub-solutions so such a training scheme will probably lead to local optima. Moreover, as another drawback of TAM and GLOP, they decompose CVRP and PCTSP into sub-TSP for the pre-trained conquering policy, but such solving framework cannot convert sub-optimal initial partitions into optimal ones, thus affecting the accessibility of optimal solutions.
@@
H-TSP [20] focuses on the TSP problem and designs a lightweight convolutional neural network and clustering mechanism for initial solution generation. Unlike other works, it implements a joint training scheme after pre-trained conquering policies. However, Section 5.1 also demonstrates that the pre-training before a joint (i.e., unified) training scheme will also lead to local optima. Moreover, similar to LCP, it does not align the optimization objectives of the dividing policy with the overall objective function of TSP as well, which further limits its effectiveness.
Appendix B Detailed Definition
B.1 Definition of Large-scale CO problems
We refer to the definition in [67] for large-scale CO problems. Large-scale CO problems have more data, decision variables, or constraints. In this article, we have considered more decision variables and data volumes in CO problems like TSP and more constraints in CO problems like MIS (i.e., dense graphs have relatively more constraints). Most large-scale NCO works solve TSPs on scales such as 500, 1000, 2000, etc. [39, 8, 16, 13], so UDC follows this setting in testing TSP and applies it to general CO problems.
B.2 Definition of Problem & Sub-problem
TSP
@@ B.2 Definition of Problem & Sub-problem
Traveling salesman problem (TSP) is one of the most representative COPs [68]. Let $\{D=d_{j,k},j=1,\ldots,N,k=1\ldots,N\}$ be the $N\times N$ cost metrics of a $N$-nodes TSP problem (in practice, transformed from the input node-wise coordinates), where $d_{j,k}$ denotes the cost between nodes $k$ and $j$, the goal is to minimize the following objective function:
@@
where the solution $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{N})$ ($\tau=N$) is a permutation of all nodes. All the feasible solutions (i.e., in $\Omega$ of Eq. (1)) satisfy the constraint of node degree being two and containing no loop with lengths less than N.
The sub-TSP in UDC is usually considered as an open-loop TSP problem [69] (i.e., SHPP in GLOP [13]). To make the optimization target of both stages correspond (with objective function $f(\cdot)$ and $f^{\prime}(\cdot)$, respectively), with the sub-cost metrics $\{D^{\prime}=d^{\prime}_{j,k},j=1,\ldots,n,k=1\ldots,n\}$, the objective function of sub-TSP solution $\boldsymbol{s}=(s_{1},s_{2},\ldots,s_{n})$ is as follow:
$\mbox{minimize}\quad f^{\prime}(\boldsymbol{s})=\sum_{t=1}^{n-1}d^{\prime}_{s_{t},s_{t+1}}$ (8)
The feasibility set of sub-TSP adds additional selection constraints for the starting and ending points (i.e., the first and last node in the original solution segments). UDC trains the TSP on various scales for every $n$ nodes, from $N$=500 to $N$=1,000 (i.e., for easier implement, we set $n$=100, and sampling $N$ from $\{500,600,700,800,900,1,000\}$). Since the nodes in each sub-problem always come from a compact region of the original instance, so we follow the normalization approach in GLOP and conduct a coordinate transformation to coordinates of sub-problems. The specific formula is detailed in Eq. (20).
CVRP
CVRP incorporates the vehicle capacity as a restriction while considering the total distance as its objective function. Each CVRP contains a depot and several customers, on a cost matrix $\{D=d_{j,k},j=0,\ldots,N,k=0\ldots,N\}$, the CVRP can be expressed as follows:
@@
$\sum_{i\in\boldsymbol{\rho}^{j}}\delta_{i}\leq C,\quad j=1,\ldots,q,$
where $\boldsymbol{x}$ is a solution representing the complete route of vehicles and consists of $q$ sub-routes $\boldsymbol{x}=\{\boldsymbol{\rho}^{1},\boldsymbol{\rho}^{2},\ldots,\boldsymbol{\rho}^{q}\}$. Each sub-route $\boldsymbol{\rho}^{j}=(\rho^{j}_{1},\ldots,\rho^{j}_{n_{j}}),\ j\in\{1,\ldots,q\}$ starts from the depot $x_{0}$ and backs to $x_{0}$, $n_{j}$ represents the number of customer nodes in it. $n=\sum_{j=1}^{q}n_{j}$ is the total number of customer nodes; $\delta_{i}$ denotes the demand of node $i$; $C$ denotes the capacity of the vehicle. In generating CVRP instances in test sets, we follow the common setting in [13, 14], setting $C=100$ for 500-node CVRP (i.e., CVRP500), $C=200$ for CVRP1,000, $C=300$ for CVRP2,000 and larger scale. In training, the capacity $C$ is uniformly sampled from 100 to 200 [6, 16].
Similar to sub-TSP, Sub-CVRP is an open-loop CVRP (not forcing both starting and ending at depot $x_{0}$ for sub-route $\boldsymbol{\rho}^{1}$ and $\boldsymbol{\rho}^{q}$) and also restricts the selection of starting and ending nodes. Moreover, each sub-CVRP instance has a special constraint on the capacity of the first and last sub-routes to ensure the legality of the merged solution. As shown in the right half of Figure 4, the capacity of the first and last sub-route is restricted to be lower than the rest capacity of their connected partial sub-routes. Afterward, each pair of adjunct capacity constraints (shown in Figure 4) is normed to make the sum of 1. We follow the setting of BQ-NCO [14] in representing CVRP solutions, converting the route into a solution and an aligned solution flag which is set to 1 for nodes before a depot.
Figure 4: The process of preparing sub-CVRPs.
OP
OP assigns each node a prize $\rho_{i}$ and the goal is to maximize the total prize of all collected nodes, while keeping the total length of the route (a loop through all collected nodes) not surpassing a maximum length $T$. OP contains a depot node serving as the route’s starting node. We evaluate the UDC on OP500, OP1,000, and OP2,000 and set the $T$ on all scales to 4. Assume solution route of OP is an $\tau$-length ordered sequence $\boldsymbol{x}=(x_{0},\ldots,x_{\tau})$ ($x_{0}$ represents be depot, $\rho_{0}$ = 0) and $\{D=d_{j,k},j=0,\ldots,N,k=0\ldots,N\}$ is the $N\times N$ cost metrics of a $N$-nodes TSP problem, the objective function can be formulated as:
maximize $f(\boldsymbol{x})=\sum_{i\in\boldsymbol{x}}\rho_{i},$ (10)
@@
We follow the general OP (distance) (i.e., OP(dist)) setting to generate OP instances [50, 51, 61, 14], whose prize of each node is calculated $\rho_{i}$ based on its distance to the depot. In practice, we follow the setting in AM [50], $\hat{\rho}_{i}=1+\lfloor 99\frac{d_{0,i}}{\text{max}_{j=1}^{N}d_{0j}}\rfloor$, $\rho_{i}=\frac{\hat{\rho_{i}}}{100}$. The sub-OP in UDC contains a segment of the original solution while also randomly introducing some unselected nodes. By solving these sub-OPs, the conquering stage gains the potential to improve a low-quality initial solution to optimal. For sub-solution $\boldsymbol{s}$, the objective function of sub-OP is $f^{\prime}(\boldsymbol{s})=\sum_{i\in\boldsymbol{s}}\rho_{i}$. There is no normalization step in the conquering stage when solving OP. Sub-OP restricts the lengths of the generated sub-solution not surpassing the total length of the original sub-OP solution. One of the $\lfloor\frac{N}{n}\rfloor$ sub-OPs contains the depot and the depot is restricted to be included in the sub-solution of this sub-OP.
In testing, solving sub-OP limits the path length not to exceed the original sub-solution length, so continuously executing the conquering stage will result in a gradually shorter total route length $l(\boldsymbol{x}_{r})$. So, we count the margin (i.e., $T-l(\boldsymbol{x}_{r})$) of the constraint $T$ in the current solution. In each conquering stage, we add the margin $T-l(\boldsymbol{x}_{r})$ to the length constraint of a random sub-OP.
PCTSP
With a similar optimization requirement compared to OP, PCTSP is also a typical CO problem [50]. PCTSP gives each node a prize $\rho_{i}$ and a penalty $\beta_{i}$. The goal is to minimize the total length of the solution route (a loop through all collected nodes as well) plus the sum of penalties of unvisited nodes. Moreover, as a constraint, PCTSP solutions should collect enough total prizes. The objective function is defined as follows:
@@
We follow the settings in AM [50] and GLOP [13] to generate PCTSP instances, $\rho_{i}\sim\text{Uniform}(0,\frac{4}{N})$, $\beta_{i}\sim\text{Uniform}(0,3\cdot\frac{K^{N}}{N})$. We follow the setting in [13], define $K^{N}=9,12,15$ for $N=500,1,000,2,000$.
Compared to OP, UDC uses a similar way to process sub-PCTSP, introducing random unselected nodes in generated sub-problems and restricting the appearance of the depot. As a similar constraint of sub-OP, sub-PCTSP also restricts the total prize collected in the generated sub-solution to be larger than the total prize in the original sub-solution. Specifically, sub-PCTSP normalizes prizes of each node to $\rho_{i}^{\prime}$ by dividing this constraint for better alignment, and the necessity of this procedure is demonstrated in Appendix E with ablation studies. The objective function of sub-PCTSP solution $\boldsymbol{s}$ is as follows:
minimize $f^{\prime}(\boldsymbol{s})=\sum_{t=1}^{\tau-1}d_{s_{t},s_{t+1}}+\sum_{i\notin\boldsymbol{s}}\beta_{i},$ (12)
subject to $\sum_{i\in\boldsymbol{s}}\rho_{i}^{\prime}\geq 1.$
The existing method GLOP [13] divides the PCTSP into sub-TSP, this problem-specific design not only harms the accessibility to optimal solutions but also impairs the generation ability of the framework to general CO problems.
SPCTSP
SPCTSP evaluates the capability of the proposed UDC in dealing with uncertainty. As a notable difference to PCTSP, in SPCTSP, the real collected prize of each node only becomes known upon visitation but an expected node prize $\rho^{*}_{i}\in\text{Uniform}(0,2\rho_{i})$ is known upfront. In solving SPCTSP with UDC, the expected node prize $\rho^{*}_{i}$ is used to process sparse graph $\mathcal{G}_{D}$, heatmap $\mathcal{H}$, and the encodings of conquering policy, while the real prize $\rho_{i}$ is only used to process the current states and constraints when constructing solutions of SPCTSP and sub-SPCTSP.
KP
Besides VRPs, KP is also a typical CO problem which can be formulated as follows:
minimize $f(\mathbb{X})=-\sum_{i\in\mathbb{X}}v_{i},$ (13)
subject to $\sum_{i\in\mathbb{X}}w_{i}\leq W,$
where $\mathbb{X}\subseteq\{1,2,...,N\}$ represents the selection item indexes (i.e., KP solution) in a knapsack and $v_{i}\sim\text{Uniform}(0,1)$ and $w_{i}\sim\text{Uniform}(0,1)$ is the value and weight of candidate objects. In each conquering step, sub-KPs are randomly sampled. Sub-KP and KP are identical in form and objective function, with only distinction on the number of decision variables and total capacity $W$. In training, UDC disables the Reunion stage in the DCR training method for KP, and the capacity $W$ is sampled from 50 to 100. In testing, UDC adopts a similar approach to OP in introducing the margin of maximum capacity (i.e., $W-\sum_{i\in\mathbb{X}}w_{i}$) to sub-KPs, adding the current margin (i.e., capacity minus current load) to a random sub-KP.
OVRP
As a variant on CVRP, OVRP calculates the objective function of each sub-route as open-loops. The objective function of OVRP is calculated as follows:
@@
ATSP
ATSP is a variant version of TSP with asymmetric cost metrics $\{D=d_{j,k},j=0,\ldots,N,k=0\ldots,N\}$ (i.e., ATSP and sub-ATSP have the same objective function as TSP and sub-TSP, respectively). Different from TSP, ATSP instances are inputted as cost metrics directly so this specificity makes it impossible to use conquering models such as POMO [5], ICAM [16], etc. [41] When dealing with sub-ATSP, we employ a matrix-based sub-problem normalization method to normalize the longest item (i.e., edge) in cost metrics of sub-ATSPs to 1. For testing scale, the 1000-node ATSP is already huge in terms of the data size, while the 2000-node data is unable to generate and store in memory, so we refer to the testing scales of GLOP [13], conducting experiments on the 250-node (i.e., ATSP250), the ATSP500, and the ATSP1000.
MIS
MIS is originally defined on a graph $\mathcal{G}=\{\mathbb{V},\mathbb{E}\}$ where $N=|\mathbb{V}|$ so in the dividing step we have $\mathcal{G}_{D}=\mathcal{G}$ directly (without additional operations to construct the sparse graph $\mathcal{G}_{D}$). For solution set $\mathbb{X}\subseteq\{1,2,...,N\}$, the objective function of MIS is to maximize the number of visited nodes in solution (i.e., $f(\mathbb{X})=|\mathbb{X}|$) while ensuring no edge connections $\boldsymbol{e}\in\mathbb{E}$ between any two visited nodes. In training, follow the MIS instance generation process of ER[700-800] with $p$=0.15 [35]. The nodes in sub-MIS are randomly selected and nodes connecting to visited nodes are constrained as unaccessible. The objective function of a sub-MIS solution $\mathbb{S}$ is $f(\mathbb{S})=|\mathbb{S}|$. The conquering policy for sub-MIS uses a Heatmap-based solver as well (i.e., AGNN in UDC) with non-autoregressive construction methods.
Min-max mTSP
As a variant of VRP, the min-max vehicle routing problem (min-max VRP), instead of minimizing the total length of all sub-routes (i.e., min-sum), seeks to reduce the length of the longest one among all the sub-routes (i.e., min-max). Min-max VRP has important application value in real life as well [70, 71] and researchers have also paid attention to solving them with deep learning techniques [72, 63]. The solution of min-max mTSP is defined similarly to CVRP solutions, consisting of exactly $M$ sub-routes (as a constraint on the number of total sub-routes) $\boldsymbol{x}=\{\boldsymbol{\rho}^{1},\boldsymbol{\rho}^{2},\ldots,\boldsymbol{\rho}^{M}\}$. The objective function of min-max mTSP is to minimize the length of the longest sub-route, which is as follows: